A plane can fly 362 mph in still air. If it can fly 210 miles downwind in the same amount of time it can fly 152 miles upwind, find the velocity of the wind.

Answer by mananth(16949)

plane speed = 362 mph
wind speed = x mph
wind speed
speed against wind 362 - x
speed with wind 362 + x

Distance against wind 152 miles
Distance with wind 210 miles

t=d/r time against = time with wind
152/(362-x)=210 /(362+x)
152(362 +x)=210 (362-x)
55024+ 152x=76020-210x
210x+152x=20996
362x=20996
/ 362
x=58mph wind speed
m.ananth@hotmail.ca